HUBBLE SPACE TELESCOPE DAILY REPORT #2045 PERIOD COVERED: 0000Z (UTC) 01/22/98 - 0000Z (UTC) 01/23/98 Daily Status Report as of 023/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of NIC/1/2/3 7235 (Observations of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Hubble Deep Field area HDF-123645. This program observes a portion of the Hubble Deep Field with all three NICMOS cameras in the F160W and F110W filters to produce infrared images equivalent in depth to the existing WF/PC-2 HDF images. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Four Sets of STIS/MA1/MA2 7920 (STIS MAMA Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make parallel field, ultraviolet imaging to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 7900 (STIS CCD Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, observations with the NIC/3 Camera. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as scheduled, and no anomalies were reported. 1.6 Completed Three Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/3 7814 (NICMOS Camera 3 Pointed Super Flats) The Near Infrared Camera (Camera Three) was used to make a "pointed super flat" calibration image. The observations will be used to build high S/N NICMOS flat fields for the different filters. Special commanding is required to move the NICMOS FOM. The observations was completed as planned, and no anomalies were reported. 1.8 Completed NIC/3 7459 (The Age and Content of a {Proto} Galaxy Cluster at z=2.39) The Near Infrared Camera (Camera Three) was used to make observations of the proto-galaxy cluster Gal-Clust-1239. Our WF/PC-2 observations of the field around the radio galaxy 53W002 {z=2.39} have discovered a remarkable population of Lyman Alpha emission-line objects at the same redshift, with 18 identified in a single WF/PC-2 mosaic. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Operations Notes: The Battery #2 Capacity Check Test is ongoing. Discharge was terminated nominally at 15 Volts by a special hardware test limit at 022/1551Z. This yielded an estimated capacity of 74.0 A-h down to 26.37 Volts. (The previous test in April 1997 yielded 78.5 A-h.) The battery was connected to solar array section 2 at 022/1707Z. Charging started at 022/1723Z, a normal orbit day entry. The NICMOS-3 Hubble Deep Field (HDF) campaign continues. As indicated in the proposals described above, all associated observations appeared to execute nominally, and related guide star performance was also nominal. Operations personnel report that the NICMOS instrument is performing excellently. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH